# FT-Shield: A Watermark Against Unauthorized Fine-tuning in Text-to-Image Diffusion Models

### v1

@@ Abstract
Text-to-image generative models based on latent diffusion models (LDM) have demonstrated their outstanding ability in generating high-quality and high-resolution images according to language prompt. Based on these powerful latent diffusion models, various fine-tuning methods have been proposed to achieve the personalization of text-to-image diffusion models such as artistic style adaptation and human face transfer. However, the unauthorized usage of data for model personalization has emerged as a prevalent concern in relation to copyright violations. For example, a malicious user may use the fine-tuning technique to generate images which mimic the style of a painter without his/her permission. In light of this concern, we have proposed FT-Shield, a watermarking approach specifically designed for the fine-tuning of text-to-image diffusion models to aid in detecting instances of infringement. We develop a novel algorithm for the generation of the watermark to ensure that the watermark on the training images can be quickly and accurately transferred to the generated images of text-to-image diffusion models. A watermark will be detected on an image by a binary watermark detector if the image is generated by a model that has been fine-tuned using the protected watermarked images. Comprehensive experiments were conducted to validate the effectiveness of FT-Shield.
@@ 1 Introduction
Generative models, particularly Generative Diffusion Models (GDMs) (Ho et al., 2020; Song et al., 2020b; Ho & Salimans, 2022; Song et al., 2020a), have witnessed significant progress in generating high-quality images from random noise. Lately, text-to-image generative models leveraging latent diffusion (Rombach et al., 2022) have showcased remarkable proficiency in producing specific, detailed images from human language descriptions. Based on this advancement, fine-tuning techniques such as DreamBooth (Ruiz et al., 2023) and Textual Inversion (Gal et al., 2022) have been developed. These methods enable the personalization of text-to-image diffusion models, allowing them to adapt to distinct artistic styles or specific subjects. For instance, with a few paintings from an artist, a model can be fine-tuned to adapt to the artistic style of the artist and create paintings which mimic the style. However, the proliferation of these methods has sparked significant copyright concerns. There are particular concerns about the potential misuse of these techniques for unauthorized style imitation or the creation of deceptive human facial images. Such actions infringe on the rights of original creators and breach the sanctity of intellectual property and privacy.
Figure 1: Generated images from fine-funed text-to-image models: a comparison of watermarking methods at different fine-tuning steps.
To protect data from being unauthorizedly used for text-to-image model fine-tuning, there is a line of research (Van Le et al., 2023; Liang et al., 2023; Shan et al., 2023; Salman et al., 2023) focusing on designing perturbations in the data to prevent any model learning from the data. In these works, adversarial techniques are applied to make the images resistant to subject-driven synthesis techniques, causing the text-to-image model to learn subjects that are substantially divergent from the original intentions. However, one issue is that these protective methods can inadvertently disrupt authorized uses of the safeguarded images. In reality, we may still allow some people to learn the image (such as for academic research purpose). To tackle this challenge, another line of research (Ma et al., 2023; Wang et al., 2023) considering watermarking techniques. Unlike the previous methods, watermarking does not disrupt the regular fine-tuning process of the models. It allows the intellectual property to be used for proper reasons, while also acting as a warning and a way to collect proof against improper uses. Specifically, watermarking techniques work via embedding a covert signal or pattern into the original data which is imperceptible to human eyes but can be detected by a specialized watermark detector. When the watermarked images are used for fine-tuning text-to-image models, the resulting generated images are expected to inherit this watermark, acting as an indelible signature. By employing a watermark detector, one can identify the presence of these watermarks in the generated content. Thus, if an entity produces images using the protected data, the watermark will serve as an evidence of infringement.
There have been existing watermarking methods (Ma et al., 2023; Wang et al., 2023) designed for protecting the copyright of images against text-to-image model fine-tuning. However, they often require sufficient fine-tuning steps to ensure that the watermark on the images can be adequately assimilated by the generative model and subsequently, transferred to the generated images. In real practice, even when the fine-tuning steps are not extensive enough for the watermarks to be learned, the style of the images can already be acquired by the model. As shown in Figure 1, as the fine-tuning steps increase, the style of the paintings can be learned by the model earlier than the watermark by Ma et al. (2023). Therefore, the watermarks may fail to act as an effective protection because the offender can evade the watermark by simply reducing the fine-tuning steps. One key reason for this issue is that those watermark generating procedures predominantly adhere to traditional watermarking strategies which are intended to trace the source of an image rather than protect the copyright against diffusion models. Therefore, integral features of their watermarks may not be promptly and precisely assimilated by the diffusion models.
@@
In this work, we propose a novel watermarking framework, FT-Shield, tailored for data’s copyright protection against the Fine-Tuning of text-to-image diffusion models. In particular, we introduce a training objective which incorporates the training loss of the diffusion model for the generation of the watermark. By minimizing this training objective, we ensure that the watermark can be quickly learned by the diffusion model in the very early stage of fine-tuning. As shown in Figure 1, by using our method, even when the style has not been learned by the diffusion model, the watermark can already be learned by the diffusion model and detected on the generated images. The effectiveness of the proposed method is verified through experiments across various fine-tuning techniques including DreamBooth, Textual inversion, Text-To-Image Fine-tuning (von Platen et al., 2022) and LoRA (Hu et al., 2021), applied to both style transfer and object transfer tasks across multiple datasets.
@@ 2.1 Text-to-image diffusion model and their fine-tuning methods
Diffusion models (Ho et al., 2020; Song et al., 2020b; Ho & Salimans, 2022; Song et al., 2020a) have recently achieved remarkable advancements in the realm of image synthesis, notably after the introduction of the Denoising Diffusion Probabilistic Model (DDPM) by (Ho et al., 2020). Building upon the DDPM framework, Rombach et al. (2022) presented the Latent Diffusion Model (LDM). Unlike conventional models, LDM conducts the diffusion process within a latent space derived from a pre-trained autoencoder, and generates hidden vectors by diffusion process instead of directly generating the image in pixel space. This strategy enables the diffusion model to leverage the robust semantic features and visual patterns imbibed by the encoder. Consequently, LDM has set new standards in both high-resolution image synthesis and text-to-image generation.
Based on the advancement in the text-to-image diffusion model, multiple fine-tuning techniques  (Gal et al., 2022; Ruiz et al., 2023; Hu et al., 2021) have been developed. These methods enable the personalization of text-to-image diffusion models, allowing them to adapt to distinct artistic styles or specific subjects. Specifically, DreamBooth (Gal et al., 2022) works by fine-tuning the denoising network of the diffusion model to make the model associate a less frequently used word-embedding with a specific subject. Textual Inversion (Ruiz et al., 2023) tries to add a new token which is bound with the new concept to the text-embedding of the model. And LoRA (Hu et al., 2021) proposed to add pairs of rank-decomposition matrices to the existing weights of the denoise network and only train the newly added weights in the fine-tuning process.
@@ 2.2 Image protection methods
In literature, there are two main types of protection methods for intellectual property against text-to-image model fine-tuning: (1) designing perturbations in the data to prevent any model learning from the data. (2) designing watermarks so that protectors can detect whether generated images infringe the intellectual property or not.
Adversarial methods. Adversarial methods protect the data copyright by taking the advantage of the idea in evasion attacks. They treat the unauthorized generative models as the target model of an attack, and the attack tries to reduce generative models’ ability in learning from images. Thus, the protected images are designed as adversarial examples which reduce the unauthorized models’ performance and further protect themselves. GLAZE is first proposed to attack the extracted feature by the encoder part in Stable Diffusion and prevent the style of images from being learned by others (Shan et al., 2023). Van Le et al. (2023) proposed Anti-DreamBooth to prevent a method of fine-tuning Stable Diffusion, DreamBooth (Ruiz et al., 2022). Liang et al. (2023) use a Monte Carlo method to generate adversarial examples to evade the infringement from Textual Inversion (Gal et al., 2022). (Salman et al., 2023) propose to modify the pictures to protect them from the image editing applications by Stable Diffusion in case the pictures are used to generate any images with illegal or abnormal scenarios. Although these methods can protect the data copyright, it destroys the authorized uses for certain generation purpose. Thus, tracing the usage and identifying the unauthorized infringement are important.
Watermarking methods. Watermarking techniques have also been considered to protect the intellectual property of images against text-to-image model fine-tuning.  Wang et al. (2023) proposed to apply an existing backdoor method  (Nguyen & Tran, 2021) to embed unique signatures into the protected images. This aims to inject extra memorization into the text-to-image diffusion models fine-tuned on the protected dataset so that unauthorized data usage can be detected by checking whether the extra memorization exists in the suspected model. However, a limitation of their approach is the assumption that the suspicious diffusion model is readily accessible to the data protector. This might not be realistic, as malicious entities might only disclose a handful of generated images but hide the fine-tuned model. Another work by  Ma et al. (2023) introduces a method that trains a watermark generator and detector simultaneously. The detector is then further fine-tuned using the images generated by the fine-tuned model. However, as discussed in Section 1, the issue of this technique is that it provides no guarantee that their watermark can be learned by the model earlier than the image’s style. The shortcomings of current methods have highlighted the need for a more robust and effective watermarking scheme.
@@ 3 Method
In this section, we first define the problem and introduce the necessary notations. Then we present the details of the proposed FT-Shield by introducing the watermark generation procedure and the training method for the watermark detector.
@@ 3.1 Problem formulation
In the scenario of copyright infringement and protection considered in this work, there are two roles: (1) a data protector that possesses the data copyright, utilizes watermarking techniques before the data are released, and tries to detect if a suspected model is fine-tuned on the protected images, and (2) a data offender that utilizes the protected data for text-to-image diffusion model fine-tuning without the permission from the data protector. The data offenders have complete control over the fine-tuning and sampling processes of the text-to-image diffusion models, while the data protectors can only modify the data they own before the data are released and access images generated by the suspected model.
As shown in Figure 2, the protection process consists of two stages: the protection stage and the audit stage. In the protection stage, the data protector protects the images by adding imperceptable watermarks to the images. Specifically, given that the size of the protected dataset is $N$, the target is to generate sample-wise watermark $\delta_{i}$ for each protected image $x_{i},\forall i=1...N$. Then these watermarks are embedded into the protected images $\hat{x}_{i}=x_{i}+\delta_{i}$. Correspondingly, the data protector will train a binary watermark detector, $D_{w}$, to test whether there is a watermark on the suspect image. To ensure that the watermarks will not lead to severe influence of the image quality, we limit the budget of the watermark by constraining its $l_{\infty}$ norm ($\|\delta_{i}\|_{\infty}\leq\eta$) to control the pixel-wise difference between the two images $x_{i}$ and $\hat{x}_{i}$. In the subsequent audit stage, if the protectors encounters suspected images potentially produced through unauthorized text-to-image models fine-tuning, they utilize the watermark detector $D_{\omega}$ to ascertain whether these images have infringed upon their data rights.
@@ 3.2 Watermark generation
As discussed in Section 1, the challenge of building a robust and reliable watermark framework is to ensure that the watermark can be quickly and precisely integrated to the diffusion models. This integration should occur as rapidly as the model learns the style during the incremental fine-tuning steps. In essence, the pivotal aim is to generate specific features that are highly sensitive to the diffusion model, ensuring the model prioritizes learning the watermark features along with or before the style of the images. To achieve this goal, given $n$ samples to be protected, we construct a training objective for the watermark as follows:
$\min_{\{\delta_{i}\}_{i\in[n]}}\min_{\theta_{1}}\sum_{i\in[n]}L_{dm}(\theta_{1},\theta_{2},x_{i}+\delta_{i},c)\text{ s.t. }\|\delta_{i}\|_{\infty}\leq\eta\vspace{-0.1cm}$ (1)
@@
where $\theta_{1}$ represents the parameters of the UNet (Ronneberger et al., 2015), which is the denoise model within the text-to-image model structure, $\theta_{2}$ denotes the parameters of the other part of the diffusion model, and $c$ is the prompt for the image generation. The function ${L}_{dm}$ denotes the fine-tuning loss of the text-to-image diffusion model:
${L}_{dm}\left(\theta_{1},\theta_{2},x_{0},c\right)=\mathbb{E}_{x_{0},t,c,\epsilon\in N(0,I_{d})}\left\|\epsilon-\epsilon_{\theta_{1},\theta_{2}}(x_{t},t,c)\right\|_{2},$ (2)
with $d$ as the dimension of the training images in the latent space, $x_{0}$ as the input image, $t$ as the timestep and $x_{t}$ as the input image with $t$ steps’ noise in the diffusion process. The above training objective aims to figure out the best perturbation $\delta_{i}$ for each sample $x_{i}$ so that the loss of a diffusion model trained on these perturbed samples can be minimized. In other words, we want to ensure that with the existence of the optimized watermark, the training loss of the diffusion model can be reduced rapidly in its training process. Intuitively, this process engenders a “shortcut” feature within the watermark which can be quickly learned and emphasized by the diffusion model. By ensuring that the model rapidly assimilates this shortcut feature, we enhance the probability that the watermark can be successfully transferred to images generated by the model fine-tuned on the watermarked datasets, even when the fine-tuning does not involve many steps. We solve this bi-level optimization problem by alternatively updating the perturbation and the parameters of the diffusion model. Details about the algorithm are provided in Appendix A.
@@
The Design of Fine-tuning Loss and Prompt. There have been multiple fine-tuning methods for text-to-image diffusion model and each of them involves different formulations of fine-tuning loss and different parts of the model to be updated. Therefore, it is crucial to ensure that our watermarking technique remains effective across different implementations of text-to-image model fine-tuning. This requires us to carefully select the specific $L_{dm}$ in Equation 1 and the caption $c$ used in watermark training. In terms of the particular $L_{dm}$ considered in Equation 1, to maintain simplicity and coherence, we focused on the most fundamental Text-to-Image Fine-tuning Method (von Platen et al., 2022). This method aligns with the original training objectives of the text-to-image model and involves updates only to the UNet. The experiments in Section 4.2 demonstrate that the watermarks generated with this formulation are able to be assimilated well by different fine-tuning methods, even those do not modify the UNet of the model such as Textual Inversion (Ruiz et al., 2023). For the caption $c$, we employ a simplistic and consistent caption format for every image within the dataset in the watermark generation process. This consistency ensures robustness in varying conditions. Specifically, for the images associated with style transfer tasks, the caption is “A painting by *”, with * denoting the artist’s name. For images used for object transfer, each is labeled by “a photo of *”, where * indicates the object’s category, such as ‘dog’, ‘toy’, or ‘person’.
3.3 Watermark detector
After the watermarks are generated, we require a binary watermark detector to detect if an image generated by the suspected model contains a watermark or not. However, considering that the size of the protected datasets is small, directly using the watermarked images and clean images to train the detector would result in poor performance due to overfitting. To tackle this challenge, we propose to use generated images from the fine-tuned text-to-image models to augment the dataset for training the detector. In the following, we first introduce the training procedure of the detector and then give the details about how to augment the training data of the detector.
Detector Training. We transform the detection of watermark into a binary classification task distinguishing clean and watermarked image. The training objective of the detector, i.e. the binary classifier $D_{\omega}$, is formulated as:
$L_{cl}\left(\omega\right)=\sum\nolimits_{i\in[n]}\left[-\operatorname{log}(1-D_{\omega}(x_{i}))-\log D_{\omega}(x_{i}+\delta_{i}\right)],$ (3)
which is the cross-entropy loss of the classification for determining whether an images is embedded with a watermark or not, and $\omega$ refers to the parameters of $D$. To save memory and computation cost, we consider applying the classification directly to the latent features of each image. In other words, we need to apply a VAE within the latent diffusion model’s structure to encode each image into its latent space before they are used to train the classifier.

### v2

@@ Abstract
Text-to-image generative models, especially those based on latent diffusion models (LDMs), have demonstrated outstanding ability in generating high-quality and high-resolution images from textual prompts. With this advancement, various fine-tuning methods have been developed to personalize text-to-image models for specific applications such as artistic style adaptation and human face transfer. However, such advancements have raised copyright concerns, especially when the data are used for personalization without authorization. For example, a malicious user can employ fine-tuning techniques to replicate the style of an artist without consent. In light of this concern, we propose FT-Shield, a watermarking solution tailored for the fine-tuning of text-to-image diffusion models. FT-Shield addresses copyright protection challenges by designing new watermark generation and detection strategies. In particular, it introduces an innovative algorithm for watermark generation. It ensures the seamless transfer of watermarks from training images to generated outputs, facilitating the identification of copyrighted material use. To tackle the variability in fine-tuning methods and their impact on watermark detection, FT-Shield integrates a Mixture of Experts (MoE) approach for watermark detection. Comprehensive experiments validate the effectiveness of our proposed FT-Shield.
@@ 1 Introduction
Generative models, particularly Generative Diffusion Models (GDMs) (Ho et al., 2020; Song et al., 2020b; Ho & Salimans, 2022; Song et al., 2020a), have witnessed significant progress in generating high-quality images from random noise. Recently, text-to-image generative models leveraging latent diffusion (Rombach et al., 2022) have showcased remarkable proficiency in producing specific, detailed images from human language descriptions. Based on this advancement, fine-tuning techniques such as DreamBooth (Ruiz et al., 2023) and Textual Inversion (Gal et al., 2022) have been developed. These methods enable the personalization of text-to-image diffusion models, allowing them to adapt to distinct artistic styles or specific subjects. For instance, with a few paintings from an artist, a model can be fine-tuned to adapt to the artistic style of the artist and create paintings which mimic the style. However, the proliferation of these methods has sparked significant concerns about the potential misuse of these techniques for unauthorized style imitation or the creation of deceptive human facial images. Such actions can potentially violate creators’ rights and compromise intellectual property (IP) and privacy integrity (Chen et al., 2023; Wang et al., 2024b; Wen et al., 2023).
Watermarking has emerged as a popular technique for protecting data’s IP against various forms of infringement (Cox et al., 2002; Podilchuk & Delp, 2001; Navas et al., 2008; Zhu et al., 2018). It works by injecting imperceptible signals or patterns into images which can later be identified by a watermark detector. This enables the tracking of unauthorized copies and facilitates the assertion of copyright infringement (Cox et al., 2002; Podilchuk & Delp, 2001; Navas et al., 2008; Zhu et al., 2018; Ren et al., 2024). Compared to other protection methods such as encryption (Forouzan, 2007), watermarking enjoys several advantages such as its stealthy nature, resilience to manipulation, and the ability to trace and manage digital assets effectively. Given its advantages, the watermarking technique also has the potential for IP protection for text-to-image model fine-tuning. When the watermarked images are used for fine-tuning text-to-image models, the resulting generated images are expected to inherit the watermark, acting as an indelible signature. By employing the detector, the presence of the watermark in the generated content can be identified, providing evidence of IP infringement.
However, we face tremendous challenges in developing watermarking techniques to prevent unauthorized fine-tuning. First, according to prior theoretical findings (Allen-Zhu & Li, 2022), the fine-tuning process of neural networks involves a sequential pattern in feature learning: certain features can be learned earlier while others are acquired later. This indicates that designing what features to embed within the watermark is crucial. If the watermark’s pattern cannot be assimilated by the model prior to style-related features, the infringer can easily circumvent the watermarks by reducing the fine-tuning steps, allowing the model to adopt the style without acquiring the watermark. Although there are recent watermarking methods proposed for images’ IP protection against text-to-image model fine-tuning, this challenge has not been solved. As shown in Figure 1, as the fine-tuning steps increase, the watermark proposed by Ma et al. (2023) is acquired by the generative model later than the targeted style, indicating an ineffective protection provided by the watermark. One potential reason is that the watermark-generating procedure predominantly adheres to traditional watermark strategies, which are intended to trace the source of an image rather than protecting its IP in the context of diffusion models fine-tuning. Consequently, there’s no assurance that the watermark’s features will be learned by the model before other features.
Second, as indicated by Ma et al. (2023), due to the distribution shift between the fine-tuning data of generative models and the resulting generated images, it is crucial to incorporate images generated by fine-tuned models to develop the watermark detector. Since there are numerous fine-tuning methods introduced from different perspectives, a detector for one method might lose its effectiveness for others. This issue is highlighted by Ma et al. (2023) that a large drop in watermark detection performance is observed when applying a detector tailored to one fine-tuning method to others. Meanwhile, in reality, a data protector may not know which fine-tuning method was used by the infringer, thus it is desired to design a watermark detection strategy that is effective for various fine-tuning methods.
Figure 1: An illustration of generated images from fine-tuned text-to-image models w.r.t. fine-tuning steps. While previous work (Ma et al., 2023) requires extensive fine-tuning to ensure that the watermark is learned, our method enables the watermark to be learned in the early stages of fine-tuning. Prompt used for the generation: Cherry blossoms in full bloom. Targeted style: the style of artist Beihong Xu.
@@
To tackle the aforementioned challenges, we propose a novel watermarking framework, FT-Shield, tailored for data’s copyright protection against the Fine-Tuning of text-to-image diffusion models. In particular, we introduce a training objective incorporating the fine-tuning loss of diffusion models for watermark generation. By minimizing the objective, we ensure that the optimized watermark pattern can be quickly learned by diffusion model at the very early stage of fine-tuning. As shown in Figure 1, even when the style has not been adopted, our watermark has already been learned by the diffusion model. Furthermore, to obtain a watermark detector for various fine-tuning methods, we introduce a detection framework based on Mixture of Experts (Jacobs et al., 1991). The effectiveness of FT-Shield is verified through experiments across various fine-tuning approaches including DreamBooth (Ruiz et al., 2023), Textual Inversion (Gal et al., 2022), Text-to-Image Fine-Tuning (von Platen et al., 2022) and LoRA (Hu et al., 2021), applied to both style transfer and object transfer tasks across multiple datasets.
@@ 2.1 Text-to-image diffusion model and their fine-tuning methods
Diffusion models (Ho et al., 2020; Song et al., 2020b; Ho & Salimans, 2022; Song et al., 2020a; Dhariwal & Nichol, 2021) have recently achieved remarkable advancements in the realm of image synthesis, notably after the introduction of the Denoising Diffusion Probabilistic Model (DDPM) by Ho et al. (2020). Building upon the DDPM framework, Rombach et al. (2022) presented the Latent Diffusion Model (LDM). Unlike conventional models, LDM conducts the diffusion process within a latent space derived from a pre-trained autoencoder, and generates hidden vectors by diffusion process instead of directly generating the image in pixel space. This strategy enables the diffusion model to leverage the robust semantic features and visual patterns imbibed by the encoder. Consequently, LDM has set new standards in both high-resolution image synthesis and text-to-image generation. Building upon text-to-image diffusion models, multiple fine-tuning techniques (Gal et al., 2022; Ruiz et al., 2023; Hu et al., 2021) have been developed. These methods enable the personalization of text-to-image models, allowing them to adapt to distinct artistic styles or specific subjects. Specifically, DreamBooth (Gal et al., 2022) works by fine-tuning the denoising network of the diffusion model to make the model associate a less frequently used word-embedding with a specific subject. Textual Inversion (Ruiz et al., 2023) tries to add a new token which is bound with the new concept to the text-embedding of the model. LoRA (Hu et al., 2021) adds pairs of rank-decomposition matrices to the existing weights of the denoise network and only trains the newly added weights in fine-tuning.
@@ 2.2 Image protection methods
To protect images’ IP from unauthorized learning by text-to-image models, in literature, two predominant methods are employed: (1) Adversarial methods which design perturbations in the data to prevent any model learning from the data; and (2) Watermarking techniques which introduce imperceptible signals to the image to enable protectors to detect infringement.
Adversarial methods. Adversarial methods protect data’s IP by applying the idea of evasion attacks. They treat the unauthorized generative models as the target of attack, and develop adversarial examples to disrupt the learning process of unauthorized fine-tuning. GLAZE (Shan et al., 2023) is the first adversarial method which focuses on attacking the features extracted by the encoder in Stable Diffusion to prevent the learning of image styles. The work of Van Le et al. (2023) and Liang et al. (2023) introduces methods to generate adversarial examples to evade the infringement from DreamBooth (Ruiz et al., 2023) and Textual Inversion (Gal et al., 2022), respectively. Additionally, Salman et al. (2023) proposed to alter the pictures to protect them from image editing applications by Stable Diffusion in case the pictures are used to generate images with illegal or abnormal scenarios. Although these methods provide effective protection against infringement, they can inadvertently disrupt authorized uses (such as for academic research purposes) of the safeguarded images. This indicates the necessity of developing watermarking approaches, which allow the IP to be used for proper reasons while also acting as a way to collect proof against improper uses.
Watermarking methods. Watermark has also been considered to protect the IP of images against unauthorized usage during the fine-tuning of text-to-image models. Wang et al. (2024a) proposed to apply an existing backdoor method (Nguyen & Tran, 2021) to embed unique signatures into the protected images. It aims to inject extra memorization into the text-to-image models fine-tuned on the protected dataset so that unauthorized data usage can be detected by checking whether the extra memorization exists in the suspected model. However, one limitation is the assumption that the suspicious model is readily accessible to the data protector. This might not be realistic, as malicious entities can hide the fine-tuned model and only disclose a handful of generated images. Another work by Ma et al. (2023) introduces a method that simultaneously trains a watermark generator and detector. The detector is then further fine-tuned with the images generated by the fine-tuned model. However, as discussed in Section 1, the issue of this technique is that it provides no guarantee that their watermark can be learned by the model earlier than the image’s style. The shortcomings of current methods highlight the need for a more robust and effective watermarking design.
@@ 3 Method
In this section, we first define the problem and introduce the key notations. Then we elaborate on some prior theoretical insights to present the challenge in watermark generation and introduce details of the watermark generation process employed by FT-Shield. Lastly, we introduce the MoE framework to detect watermarks on images generated by various fine-tuning methods.
@@ 3.1 Problem formulation
In the scenario of copyright infringement and protection considered in this work, there are two roles: (1) a data protector that possesses the data copyright, utilizes watermarking techniques before the data is released, and tries to detect if a suspected image is generated by a model fine-tuned on the protected images, and (2) a data offender that uses the protected data for text-to-image model fine-tuning without permission from the data protector. The data offenders have complete control over the fine-tuning and sampling processes of the text-to-image diffusion models, while the data protectors can only modify the data they own before their data is released and access images generated by the suspected model.
As shown in Figure 2, the protection process consists of two stages: the protection stage and the audit stage. In the protection stage, the data protector protects the images by adding imperceptible watermarks to the images. Specifically, given that the size of the protected dataset is $N$, the target is to generate sample-wise watermark $\delta_{i}$ for each protected image $x_{i},\forall i=1...N$. Then these watermarks are embedded into the protected images $\hat{x}_{i}=x_{i}+\delta_{i}$. Correspondingly, the data protector develops a watermark detection approach, denoted by function $D_{w}(\cdot)$, to test whether there is a watermark on the suspected image. To ensure that the watermarks will not lead to severe influence on image quality, we limit the budget of the watermark by constraining its $l_{\infty}$ norm ($\|\delta_{i}\|_{\infty}\leq\eta$) to control the pixel-wise difference between the two images $x_{i}$ and $\hat{x}_{i}$. In the audit stage, if the protectors encounter suspected images potentially produced through unauthorized text-to-image models fine-tuning, they will apply the watermark detection process $D_{w}(\cdot)$ to ascertain whether these images have infringed upon their data rights.
@@ 3.2 Watermark generation
As mentioned in Section 1, to ensure robust IP protection, it is crucial for the watermarks to be learned earlier in the fine-tuning process. In this subsection, we first introduce the key challenge to achieve this goal by revising how neural networks learn features in pre-training and fine-tuning. Then we propose our watermark generation approach, explaining how it effectively mitigates this challenge.
A key challenge for watermark generation. Based on the theoretical analysis on Allen-Zhu & Li (2022), due to the random initialization, each hidden node in neural networks randomly captures some features at the beginning of training. During pre-training, instead of learning other features from the data, the network emphasizes and strengthens those features that were captured at initialization and are also present in the dataset. Meanwhile, it eliminates features that, despite being learned at initialization, do not find a match in the dataset. Essentially, a well-pre-trained model captures only the features that appear in the pre-training data. An important implication of Allen-Zhu & Li (2022) is that, when learning from the fine-tuning data, the neural network can easily adapt the features that already exist in the pre-training data, but it is difficult to learn new features that only appear in the fine-tuning data. This suggests a challenge: if the watermark’s features are new to the diffusion model, it is hard to ensure that the watermark can be easily learned by the model during fine-tuning.
The proposed watermark generation approach. To overcome the above challenge, we propose to simultaneously train the watermark and fine-tune the model as follows. Given $N$ samples to be protected, we construct a training objective for the watermark as:
$\min_{\theta_{1}}\min_{\{\delta_{i}\}_{i\in[N]}}\sum_{i\in[N]}L_{LDM}(\theta_{1},\theta_{2},x_{i}+\delta_{i},c)\text{ s.t. }\|\delta_{i}\|_{\infty}\leq\eta$ (1)
@@
where $\theta_{1}$ represents the parameters of the UNet (Ronneberger et al., 2015), which is the denoise model within the text-to-image model structure, $\theta_{2}$ denotes the parameters of the other part of the diffusion model, and $c$ is the prompt for the image generation. The function ${L}_{dm}$ indicates fine-tuning loss of the text-to-image diffusion model:
$L_{LDM}(\theta_{1},\theta_{2},x_{0},c)=\mathbb{E}_{t,\epsilon\sim\mathcal{N}(0,I_{d})}\left\|\epsilon-\epsilon_{\theta_{1},\theta_{2}}(x_{t},t,c)\right\|_{2},$ (2)
with $d$ as the dimension of the training images in the latent space, $x_{0}$ as the input image, $t$ as the timestep and $x_{t}$ as the input image with $t$ steps’ noise in the diffusion process. The above training objective aims to identify the best perturbation $\delta_{i}$ for each sample $x_{i}$ so that the loss of a diffusion model trained on these perturbed samples can be minimized.
To explain the above design in Equation 1, since different features are learned differently in the fine-tuning stage, we use a minimization to find the most proper features for the watermark. The inner minimization of $\delta_{i}$ ensures that the watermark can be easily captured by the fine-tuned model, even when the fine-tuning does not involve many steps. Numerically, we solve this bi-level optimization problem by alternatively updating the perturbation and the parameters of the diffusion model. Details about the algorithm are provided in Appendix A.
@@
The design of fine-tuning loss and prompt. There have been multiple fine-tuning methods for text-to-image diffusion models and each of them involves different fine-tuning loss and different parts of the model to be updated. Therefore, it is crucial to ensure that our watermark remains effective across different types of fine-tuning. This requires a careful selection of the specific $L_{LDM}$ in Equation 1 and the caption $c$ used in watermark training. In terms of the particular $L_{LDM}$ in Equation 1, to maintain simplicity and coherence, we focused on the most fundamental Text-to-Image Fine-Tuning Method (von Platen et al., 2022). This method aligns with the original training objectives of the text-to-image model and involves updates only to the UNet. Our experiments in Section 4.2 demonstrate that the watermarks can be assimilated well by different fine-tuning methods, even for those which did not modify the UNet such as Textual Inversion (Ruiz et al., 2023). For the caption $c$, we employ a simplistic and consistent caption format for every image in the dataset in watermark generation. This consistency ensures robustness in varying conditions. Specifically, for the images associated with style transfer tasks, the caption is ‘A painting by *’ for art datasets, with * denoting the artist’s name, and ‘A Pokemon character’ for the Pokemon dataset. For images used for object transfer, each is labeled by ‘a photo of *’, where * indicates the object’s category, such as ‘toy’, or ‘person’.
3.3 Mixture of Watermark Detectors
Figure 3: An illustration of mixture of watermark detectors.
After the watermarks are generated, we need to detect if a suspected image contains a watermark or not. As mentioned in Section 1, a challenge of the watermark detection lies in the problem that the watermark detector tailored to one fine-tuning method cannot transfer well to other methods. This is because different fine-tuning methods modify various parts of the model, resulting in distinct impacts on the generation process and feature updates. To address this issue, we propose a strategy based on the Mixture of Experts (MoE) (Jacobs et al., 1991).
The general workflow. An overview of the watermark detection process using MoE is presented in Figure 3. The MoE framework comprises two main elements: the Gating Model and multiple expert models. In the watermark detection phase, the Gating Model is applied to the suspected image to estimate the likelihood that the image was produced by each potential fine-tuning method. Meanwhile, multiple experts of watermark detector, each customized for a distinct fine-tuning method, are employed. The final prediction on the presence of a watermark is then derived by taking a weighted average of these expert assessments, with the weights determined by the Gating Model’s probabilities. Formally, the MOE detection framework can be formulated as:
$D_{w}(x)=\sum_{i=1}^{M}\text{softmax}_{i}(G(x))\cdot E_{i}(x),$
where $x$ refers to the image to evaluate, $M$ refers to the total number of expert models, $G(\cdot)$ denotes the Gating Model, $E_{i}(\cdot)$ represents the i-th expert model, and $\text{softmax}_{i}(\cdot)$ is the i-th output of the softmax function.